OpenStudy (anonymous):

Remove the largest possible common factor 28x^4+14x^2+21x

OpenStudy (dumbcow):

Find greatest common factor of 28, 14, 21 28=2*14=2*2*7 14 = 2*7 21 = 3*7

OpenStudy (anonymous):

and?

OpenStudy (dumbcow):

once you find the largest factor you pull it out and divide into each of the three terms we know 7 is the largest common factor ->7(28/7 x^4 + 14/7x^2 +21/7x)

OpenStudy (dumbcow):

now look at how many x's you can pull out find the x with smallest exponent

OpenStudy (anonymous):

and?

OpenStudy (dumbcow):

and what? how many x's can we factor out?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

whats the answer?

OpenStudy (dumbcow):

im here to help you need to figure it out